$12.4divided by(17-n)=$4.14
Answer by ptaylor(1894)

$12.4divided by(17-n)=$4.14
($12.4)/(17-n)=$4.14 Multiply both sides by (17-n)
$12.4(17-n)/(17-n)=$4.14(17-n) simplifying we get
$12.4=$70.38-$4.14n subtract $70.38 from both sides
$12.4-$70.38=-$4.14n
-$4.14n=-$57.98
n=$14.0048

Ck by using calculator
Hope this helps----ptaylor